# Improving User Mental Models of XAI Systems with an Inclusive Design Approach

### v2

@@ 7.3. Threats to validity
Every study encounters potential challenges to its validity (Wohlin et al., 2000), and ours is no exception. We acknowledge several factors that could limit the generalizability and reliability of our findings.
The first issue is the demographic homogeneity. Although we had 69 participants, most of them were students and came from the same university. This may raise concerns about the broader application of our results. Our participants may not properly reflect the diversity of potential MNK or XAI users.
Moreover, the MNK domain itself presents a potential limitation as this domain is not common in XAI and HAI research. The MNK games Team Game used might not sufficiently replicate real-world scenarios. This could potentially compromise participants’ ability to accurately reason or understand the agents’ behavior. We decided to work with an MNK prototype as people are more familiar with Tic-Tac-Toe which allowed for straightforward judgment on agents’ decisions. Also, MNK gave us the option to ask a wider range of questions related to predictions, flaws, changes, and after-action reviews.
Another concern is that our participants performed proxy tasks. They predicted which square the AI agent would move next, but their predictions did not impact the game. Such proxy tasks may not reflect how users would perform in actual decision-making tasks. As Buçinca et al. showed evaluations of XAI systems with proxy tasks did not predict the results of evaluations with actual decision-making tasks. While our proxy tasks reduced cognitive load and allowed more focus on explanations, they might have limited us from creating immersive and high-stakes scenarios. An important future research would be evaluating if GenderMag-driven fixes can improve users’ performance in actual decision-making tasks.
We did not gather click log data for explanation interactions. Instead, we relied on participants’ mentions of using the explanations as an engagement indicator. Maintaining a click log could have offered more dependable insights on explanation usage. While we have screen recordings from each session, determining the time spent specifically on explanations would be difficult and time-consuming.
Addressing such threats would require additional studies using diverse empirical methods. Those methods have to make sure the findings apply widely to various XAI systems and participant demographics.
@@ 8. Conclusion
This paper investigated if inclusivity fixes can improve users’ mental models of the AI system. Our setting was an XAI prototype where participants observed two AI agents sequentially play a modified version of Tic-Tac-Toe. We report the results from our between-subject study where the “Original” group used the original version of the prototype and the “Post-GenderMag” group used the version with the inclusivity fixes:
  • RQ1: Participants who used the Post-GenderMag version of the prototype demonstrated significantly better mental models than the Original version participants (Section 4).
  • RQ2: There was no significant difference in the two groups of participants’ prediction performance (Section 5).
  • RQ3: Inclusivity fixes made the Post-GenderMag prototype more inclusive for women without hurting men and improved equity by reducing the gender gap (Section 6).
These results indicate the impacts of inclusivity fixes in improving users’ mental models and effectively increasing equity and inclusion within XAI systems. Our study highlights the potential of problem-solving style-centric inclusive design approaches to create more accessible and user-centric XAI prototypes, benefiting diverse users’ understanding of AI.
P37-PostGM: The BTW ¡Best to Worst explanation¿ and top 5 moves score ¡available in the Post-GenderMag version only¿ helped me gain a better understanding of why the AI did the things it did.
@@
Table 11. Codes and definitions
Code Definition
@@
Agent miss-selection Participant says the agent selected the wrong square because it is not the high/highest score (not following its reasoning).
@@
Table 12. All of Team J’s Fixes, why they fixed them (problem-solving styles impacted), and the changes made for the fixes
@@
Fix-1: Blurb of Instructions ”Attitude toward technology is by process,…look for documentation..” PSS: Info, Risk. – Added instructions on getting started. – Added summary of each explanation. – Added minimize button for instructions.
Fix-2: Exact values and Exact actions ”She is not clear on what she should do…..will not take this step.” PSS: Risk, Learn, Info, SE, Motiv. – Showed exact win%, loss%, and drawfor every rectangle in the BTW, STT, and OTB explanations using tooltips. – Added tooltips that informed about the exact actions offered by the next and the prior move buttons.
Fix-3: Buttons Design and Colors ”The game info…gives the hint….needs further indications.” PSS: Motiv, Info. – Made buttons look more like buttons. – Increased color contrast between the button and the instructions background. – Used undo icon with the prior move button.
@@
Fix-14: Move History “Abi will want to spend more time, because with the prior move…what it’s telling her.” PSS: Info, SE, Risk. – Team J wanted to have the feature to go back to any move from an older game. We have not implemented it.
Fix-15: Zoom in/out “Not sure that Abi would get the information…given her information processing style. PSS: Info, SE. – Team J wanted to have a zoom-in/out feature for the explanation charts. We have not implemented it.
9 of the 15 fixes: Fix-1, 2, 4, 5, 6, 7, 8, 10, and 15 were targeted to improve the explanations.

### v3

@@ 7.3. Threats to Validity
Every study encounters potential challenges to its validity (Wohlin et al., 2000), and ours is no exception. Several factors could limit the generalizability and reliability of our findings.
The first issue is the demographic homogeneity. Although we had 69 participants, most of them were students and came from the same university. This may impact the generality of our results as our participants may not properly reflect the diversity of potential MNK or XAI users.
Moreover, the MNK domain itself presents a potential limitation. We decided to work with an MNK prototype as most people are familiar with Tic-Tac-Toe, and this familiarity gave a pre-existing basis for judging the agents’ decisions. For the same reason, MNK gave us the option to ask a wide range of questions related to predictions, flaws, changes, and after-action reviews. Still, MNK games might not be representative of other XAI applications.
Recall from Section 3.1 that Agent Blue’s noise levels differed in Game 2 from the other games. This produced less determinism, which could be viewed as participants describing multiple agents rather than one. To partially mitigate this threat, we eliminated participants’ responses from the end-of-study question “Please describe the steps in the agent’s approach or method…”, which would have encouraged them to somehow generalize all three games’ behaviors into one description. Still, the varying noise levels could have influenced participants’ understanding of the AI agent during the games.
Another concern is that our participants performed proxy tasks, i.e., they predicted which square the AI agent would move next, but their predictions did not impact the game. Such proxy tasks may not reflect how users would perform in actual decision-making tasks. In Buçinca et al.’s work, evaluations of XAI systems with proxy tasks did not predict the results of evaluations with actual decision-making tasks (Buçinca et al., 2020). We used proxy tasks to reduce cognitive load and allow more focus on explanations, but they are not the same as immersive and high-stakes scenarios.
Team Game focused on Abi-like problem-solving styles to find and fix inclusivity bugs, but did not repeat the process for Tim, because they did not have enough time available to do both. When the time is too limited to do both, the GenderMag kit (Burnett et al., 2020) recommends the “Abi-first” approach because Abi provides the strongest lens for inclusivity evaluation, as previous studies have shown (Vorvoreanu et al., 2019; Guizani et al., 2022). If Team Game had also run GenderMag sessions using Tim’s perspective, their inclusivity bug finding and fixing efforts would likely have produced more and/or different changes to the Post-GenderMag explanations.
Finally, we did not gather click log data for explanation interactions. Instead, we relied on participants’ mentions of using the explanations as an engagement indicator. Maintaining a click log could have offered additional insights on explanation usage.
Addressing such threats can only be achieved with additional studies using a variety of populations, AI agents, explanation types, and empirical methods.
@@ 8. Conclusion
“There’s an ingrained societal suspicion that intentionally supporting one group hurts another. That equity is a zero sum game.” -Angela Glover Blackwell (Blackwell, 2016)
Blackwell’s concern is a powerful observation of societal misconceptions about inclusivity and equity-focused improvement efforts. Yet, sidewalk curb cuts serve as a clear counterexample, showing that including a previously underserved group of users need not be a “compromise” that trades off benefits with those already well-served.
To investigate bringing curb cut effects to XAI, we analyzed the impacts of an AI product team’s inclusivity fixes on their XAI prototype using GenderMag, an inclusive design approach that considers users’ diverse problem-solving styles to generate actionable design directions. The results revealed that, although the AI team had focused on improving inclusivity for “Abi-like” users, their fixes led to several curb-cut effects that particularly benefited not only Abi-like users, but also brought benefits to everyone:
  • Mental model concepts scores improved: Team Game’s inclusivity fixes led to Post-GenderMag participants demonstrating significantly better conceptual mental model than the Original participants (Section 4). The differences revealed which of Team Game’s inclusivity fixes brought about these changes, the inclusivity reasoning behind each of these fixes, and the AI agent-related concepts they successfully reinforced (Section 4). These explicit ties between the inclusivity reasoning and the impacts on the entire group provides evidence of these fixes’ curb-cut effects.
  • Engaging with explanations helped: Engagement with explanations significantly predicted participants’ mental model concepts scores, and Post-GenderMag participants tended toward higher engagement than the Original participants (Sections 5 and  5.2). Since almost all of Team Game’s changes were inclusivity fixes to explanations, this further evidences the presence of curb-cut effects in Team Game’s inclusivity fixes.
  • Inclusivity improved for Abi-like participants: Team Game’s underserved users were Abi-like users, and Team Game’s inclusivity fixes were particularly effective for this group. Abi-like Post-GenderMag participants achieved significantly higher mental model concept scores than their Original counterparts (Section 6.1). Further, these gains for Abi-like users were accompanied by small (not significant) gains for Tim-like users, again consistent with curb-cut effects.
  • Inclusivity improved by gender: Finally, the inclusivity fixes brought about significantly higher mental model concepts scores for an underserved gender, namely women, while also slightly increasing men’s. The increase in inclusivity reduced the gender gap by 45% (Section 6.2).
  • A curb-fence effect with predictions: However, there was also a “curb-fence” effect (opposite of the curb-cut effect): Participants’ abilities to predict the AI agent’s next move appeared to get worse. While helpful for reinforcing conceptual mental model understanding at a declarative knowledge level, Team Game’s fixes may have also encouraged participants to overrely on the explanations, affecting their abilities to predict the agent’s next move (Section 5).
In sum, these results suggest, first, that using an inclusive design to improve AI explanations can not only improve explanations’ ability to inclusively serve users who are being left behind, but also can make AI explanations “just plain better” for everyone—in essence bringing curb-cut effects to XAI. However, they also suggest the possibility of curb-fence effects, in which changes intended to increase inclusivity of explanations can be “too effective,” potentially leading to overreliance by everyone. We hope other XAI researchers will join us in exploring the nuances of inclusive design approaches’ potential to bring curb-cut effects into explanations while guarding against curb-fence effects, to address those who XAI currently underserves, while also benefiting everyone along the way.
9. Acknowledgment
We thank all the study participants for their time and effort. We also thank Bhavika Madhwani, Jeramie Kim, Elizabeth Li, Sujay Koujalgi, Iyadunni Adenuga, Mitchell Hoesing, and Jeff Schulman for their help in this paper. This work has been supported in part by NSF #1901031 and #2042324, and by USDA-NIFA/NSF #2021-67021-35344.
@@ 9. Acknowledgment
Table 10. Concepts and definitions
Concept Definition
@@
Agent misselection Participant says the agent selected the wrong square because it is not the high/highest score (not following its reasoning).
@@
Table 11. All of Team Game’s Fixes, why they fixed them (problem-solving styles impacted), and the changes made for the fixes
@@
Fix-1: Blurb of Instructions “Attitude toward technology is by process,…look for documentation..” PSS: Info, Risk. – Added instructions on getting started. – Added summary of each explanation. – Added minimize button for instructions.
Fix-3: Buttons Design and Colors “The game info…gives the hint….needs further indications.” PSS: Motiv, Info. – Made buttons look more like buttons. – Increased color contrast between the button and the instructions background. – Used undo icon with the prior move button.
@@
Fix-14: Move History “Abi will want to spend more time, because with the prior move…what it’s telling her.” PSS: Info, SE, Risk. – Team Game wanted to have the feature to go back to any move from an older game. We have not implemented it.
Fix-15: Zoom in/out “Not sure that Abi would get the information…given her information processing style. PSS: Info, SE. – Team Game wanted to have a zoom-in/out feature for the explanation charts. We have not implemented it.